OpenStudy (anonymous):

Write a possible explicit rule for the nth term of the sequence: 25, 15, 5, -5, -15, ... Please help!

OpenStudy (anonymous):

Possible answers: a)an = 35(an-1) - 10 b)an = 35n - 10 c)an = 35 - 10n d)an = 35 - 10(an-1)